OpenStudy (anonymous):

Multiply and simplify: x squared minus 4 over x squared minus 1 times x + 1 over x + 2

OpenStudy (lgbasallote):

let me translate that for you \[\huge \frac{x^2-4}{x^2-1} \times \frac{x+1}{x+2}\] any ideas now?

OpenStudy (anonymous):